OpenStudy (anonymous):

A = 9 over 2(x + 23) Determine the value of x when A = 108.

OpenStudy (anonymous):

A = 9/[2(x+23)] ? x+23 = 9/(2A) etc.

OpenStudy (anonymous):

Ummm.......lost! @douglaswinslowcooper

OpenStudy (anonymous):

x + 23 = 9/2(108) = 9/216 = 0.042 x = 0.042-23 = -22.958 check: 108 = 9/(2(-22.958+23)) = 9/2(.042) =107 (rounding errors) check.

OpenStudy (anonymous):

Still puzzled?